OpenStudy (anonymous):

Solve for y. 3y – 12 = –30 A. –6 B. –2 C. 2 D. 32

OpenStudy (anonymous):

the answer is A.)

OpenStudy (ilovepie):

y= -6

OpenStudy (gorv):

3y-12=-30 3y=-30+12 3y=-18 y=-18/3=-6

OpenStudy (anonymous):

solve for z. z - 5 / 7 = -1 -12 -2 -42 28

OpenStudy (gorv):

z-5=-1*7=-7 z=-7+5